An application of a rational function is T=(AB)/(A+B), which gives the time T, it takes for two workers to co mplete a particular task where A+B represents the time it would take for each individual worker to complete the identical task.
Estimate how long it takes you to complete a task of your choice ( house cleaning) , mowing, etc. in a given week. Suppose that Joe is slower than you at the given task and takes twice as long as you do if you work together, how long would it take you to complete the task?
Include the type of job, the time it takes you and Joe individually to complete the job and the calculations needed to show how long it woul take to complete the job if you worked together inclued units with your answer.
I have chosen House work, which takes me aprox. 2 hrs. on the other hand my mother who is handicapped, it takes her approx. 4-5 hrs to do the same type of work. Can you show me the application for this problem using the equation above?
Thank you.

Answer by stanbon(75887)

An application of a rational function is T=(AB)/(A+B), which gives the time T, it takes for two workers to co mplete a particular task where A+B represents the time it would take for each individual worker to complete the identical task.
Estimate how long it takes you to complete a task of your choice ( house cleaning) , mowing, etc. in a given week. Suppose that Joe is slower than you at the given task and takes twice as long as you do if you work together, how long would it take you to complete the task?
Include the type of job, the time it takes you and Joe individually to complete the job and the calculations needed to show how long it woul take to complete the job if you worked together inclued units with your answer.
I have chosen House work, which takes me aprox. 2 hrs. on the other hand my mother who is handicapped, it takes her approx. 4-5 hrs to do the same type of work. Can you show me the application for this problem using the equation above?
-----------------
You take A=2 hrs
Your Mom takes B=5 hrs
----------
T = (2*5)/(2+5) = 10/7 hrs = 1 3/10 hrs to do the job together.
-------------------
Cheers,
Stan H.